OpenStudy (anonymous):

\[ \sum_{n \rightarrow 2}^\infty \frac{(-5)^n}{8^{2n}}\] I get \[ \frac{-5}{8^2}(\frac{1}{1-\frac{-5}{8^2}}) \] but it seems that it is wrong (it is the intermediary step, but I think that this where I am going wrong). I fixed the indices to n-1 and switched the sum expression with \( \frac{a(1-r^n)}{1-r} \) some help?

OpenStudy (anonymous):

Ooops the \( n \rightarrow 2 \) should be \(n=2\)

OpenStudy (anonymous):

the equation a (1 - r^n) / (1 - r) tells you that you a is the first term of the sequence, and n is the number of terms but since n->infinity, and |r| < 1, it becomes a/(1-r) so what you should have had is: [(-5)^2/(8^4)] / [1 - (-5)/8^2] = 25/4416

OpenStudy (anonymous):

How did you get the \(\frac{-5}{8^2}\) twice on top?

OpenStudy (anonymous):

that is when n = 2 (which is the first term of the sequence). The index does not tell you what position each term is at, it only tells you the value at that index. so even though the index starts at 2, the first term is, in this case, (-5)^4/8^4

OpenStudy (anonymous):

........ you are right..... cant believe i missed that...... Thanks for the help and time